# Resource-Theoretic Hierarchy of Contextuality for General Probabilistic Theories

### v1

@@ Abstract
In this work we present a hierarchy of generalized contextuality. It refines the traditional binary distinction between contextual and noncontextual theories, and facilitates their comparison based on how contextual they are. Our approach focuses on the contextuality of prepare-and-measure scenarios, described by general probabilistic theories (GPTs). To motivate the hierarchy, we define it as the resource ordering of a novel resource theory of GPT-contextuality. The building blocks of its free operations are classical systems and GPT-embeddings. The latter are simulations of one GPT by another, which preserve the operational equivalences and thus cannot generate contextuality. Noncontextual theories can be recovered as least elements in the hierarchy. We then define a new contextuality monotone, called classical excess, given by the minimal error of embedding a GPT within an infinite classical system. In addition, we show that the optimal success probability in the parity oblivious multiplexing game also defines a monotone in our resource theory. We end with a discussion of a potential interpretation of the non-free operations of the resource theory of GPT-contextuality as expressing a kind of information erasure.
@@ 1 Introduction
A crucial research question in the foundations of quantum theory is to identify those features of quantum theory that constitute a true departure from the classical worldview. Addressing this question requires one to first establish a good notion of classicality, which adequately captures the classical worldview. We believe that a good notion of classicality should satisfy the following desiderata (see also [1]): (1) it endorses a principle that defines a clear boundary between aspects that pose interpretational issues and those that do not, (2) it is of broad range of applicability, (3) it is empirically testable, and (4) its violation constitutes a resource for practical applications, in particular in quantum information processing. [^1] This last desideratum is motivated by the belief that identifying the true nonclassicality of quantum theory will ultimately provide the answer to the question about the origin of the alleged quantum computational speed-up. Motivated by these desiderata, a leading notion of classicality is generalized noncontextuality [2]. [^2] In what follows we will often omit “generalized” and just talk of “(non)contextuality”. A noncontextual theory is one that is compatible with a classical realist explanation of its operational predictions —  it admits of a noncontextual ontological model. In such a model, any two experimental procedures that the theory predicts to be operationally indistinguishable also have the same ontological representation (see Section 2 for a precise definition ).
@@
A GPT system $\mathsf{B}$ is said to be at least as contextual as a GPT system $\mathsf{A}$ is if there exists a univalent simulation of $\mathsf{A}$ by a composite of $\mathsf{B}$ and a classical GPT system.
It is the resource ordering of a resource theory with objects given by GPT systems and free operations given by univalent simulations with free access to classical systems. In Section 5.1, we discuss a possible interpretation of (non-univalent) simulations as expressing a particular kind of information erasure.
@@
As is common in resource theories, the hierarchy of contextuality is not a total order — there are GPT systems such that neither is at least as contextual as the other one. It is not even a partial order, because there exist distinct GPT systems which are equivalent. For example, this is the case for all noncontextual systems, such as the classical bit and the classical trit. Therefore, the hierarchy is given by a preorder and it cannot be fully represented by a single numerical value — the “amount of contextuality” — assigned to each GPT system. However, in order to capture certain aspects of the hierarchy it is useful to define quantities which are order-preserving assignments of a number to each resource object. These are called resource monotones.
@@
Unlike Duarte and Amaral, we use the framework of GPTs. A GPT system consists of a collection of states, effects and a probability assigned to each pair of a state and an effect. The so-called simplex embeddability condition [47] states that a GPT system is noncontextual if and only if the associated GPT system can be embedded within a classical GPT system. This makes the GPT framework a suitable one to study generalized contextuality. The simplex embeddability condition thus characterizes the qualitative divide between contextual and noncontextual GPT systems. Our hierarchy of contextuality can be seen as a refinement thereof. It makes distinctions between contextual systems based on how (and how much) they violate simplex embeddability. In practice, this is achieved employing the notion of simulations of GPT systems from other GPT systems, as defined in [32]. We dedicate Section 5.2 to the discussion of the relation with other works that study contextuality in GPTs, namely [32, 44, 47, 48, 49, 50, 51, 52].
@@
In Section 2, we recall the standard treatment of generalized noncontextuality in the framework of operational theories and ontological models. This is connected to the language of general probabilistic theories in Section 3. There, we also describe (univalent) simulations of GPT systems and the excess measure. The hierarchy of contextuality is introduced in Section 4, where we prove that the classical excess is a resource monotone (Theorem 26) and construct a monotone from the parity-oblivious-multiplexing protocol. In Section 5, we discuss a possible interpretation of the non-free operations as involving information erasure and the relation to previous works on contextuality and GPTs. We conclude with a summary of the results and an outline of possible future research directions in Section 6.
@@ 2 Operational theories, ontological models and contextuality
In this work we consider prepare-and-measure scenarios associated with a single system. An operational theory associated with a prepare-and-measure scenario is defined by a list of possible preparations, measurements and the probabilities $\mathcal{P}(k|P,M)$ of obtaining the outcome $k$ of the measurement $M$ given that the system is prepared in the preparation $P$. An ontological model of an operational theory provides a realist explanation of the operational predictions of the theory [53]. It does so by stipulating the existence of an ontic state space for each given system, denoted with $\Lambda$, which is mathematically represented by a (standard Borel) measurable set. Each point $\lambda\in\Lambda$ represents an ontic state that describes all the physical properties of the system. The ontological model associates each preparation $P$ in the operational theory with a conditional probability distribution $\mu_{P}(\lambda)\equiv\mu(\lambda|P)$ over ontic states. We call these epistemic states as they represent states of knowledge about the underlying ontic states. Each outcome $k$ given a measurement $M$, denoted by $[k|M]$, is associated with a conditional probability distribution $\xi(k|\lambda,M)$. The latter corresponds to the probability of obtaining outcome $k$ given that measurement $M$ is implemented on a system in the ontic state $\lambda$. An ontological model of an operational theory reproduces the predictions of the theory via the classical law of total probability,
@@
Similarly, two measurement outcomes $[k|M]$ and $[k^{\prime}|M^{\prime}]$ are operationally equivalent if they give the same statistics for all possible preparations: $\mathcal{P}(k|M,P)=\mathcal{P}(k^{\prime}|M^{\prime},P)$ for all preparations $P$. In this case, we write $[k|M]\simeq[k^{\prime}|M^{\prime}]$. An ontological model is measurement noncontextual if any two such measurement outcomes are represented by the same response functions:
$[k|M]\simeq[k^{\prime}|M^{\prime}]\implies\xi_{[k|M]}=\xi_{[k^{\prime}|M^{\prime}]}.$ (3)
@@
We note that the notion of generalised contextuality extends to transformations also [2]. In the present work we will be concerned only with prepare-and-measure scenarios and as such we will ignore transformation contextuality.
@@ 3.1 GPT systems
Let us now discuss the mathematical description of GPT systems [54, 55, 56, 57] and the definition of contextuality in this context.
@@ A GPT system $\mathsf{A}$ is specified by two non-empty convex sets
$\Omega_{\mathsf{A}}\subseteq V_{\mathsf{A}},\qquad E_{\mathsf{A}}\subseteq V_{\mathsf{A}}^{*},$ (4)
where $V_{\mathsf{A}}$ is a normed vector space and $V_{\mathsf{A}}^{*}$ is the topological dual of $V_{\mathsf{A}}$. The spaces $\Omega_{\mathsf{A}}$ and $E_{\mathsf{A}}$ are called the state space and effect space, respectively, while their elements are states and effects of $\mathsf{A}$. We require that $E_{\mathsf{A}}$ contains the null effect $0\in E_{\mathsf{A}}$ and the unit effect $1\in E_{\mathsf{A}}$, which are the 0 and 1 constant functionals, respectively, in $V_{\mathsf{A}}^{*}$.
[…]
$e\cdot\omega\in[0,1],$ (5)
for all effects $e$ and all states $\omega$. Finally, to avoid complications arising from degenerate systems, we require that
@@
and similarly that effects can be distinguished by states.
@@
Likewise, every classical probabilistic system is a GPT system [54, §4].
@@
[…]
@@
The countable classical system $\Delta_{\mathbb{N}}$ has associated vector space $\ell^{1}$, the Banach space of sequences whose series are absolutely convergent. The state space, also denoted by $\Delta_{\mathbb{N}}$, consists of all probability measures on $\mathbb{N}$. The effect space $\Delta_{\mathbb{N}}^{*}$ consists of arbitrary sequences in $[0,1]^{\mathbb{N}}$ and forms a convex subset of $\ell^{\infty}$, the space of bounded sequences, which is the dual of $\ell^{1}$.
Besides individual systems, we will occasionally also need to refer to composite ones. In general, the composite of two GPT systems is not unique – see [58, Section 5] for an in-depth discussion of tensor products of GPT systems. However, among all the possible choices, there is a ‘minimal composite’ of two GPT systems $\mathsf{A}$ and $\mathsf{B}$, whose state and effect space merely contain the separable ones. Any meaningful composite of two GPT systems necessarily contains their minimal composite as a subsystem.
@@
Given two GPT systems $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ and $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}},V_{\mathsf{B}})$ we define the minimal composite system $\mathsf{A}\otimes\mathsf{B}$ to be given by the separable states and separable effects on $\mathsf{A}$ and $\mathsf{B}$. Formally $\mathsf{A}\otimes\mathsf{B}:=(\Omega_{\mathsf{A}\otimes\mathsf{B}},E_{\mathsf{A}\otimes\mathsf{B}},V_{\mathsf{A}\otimes\mathsf{B}})$ where,
@@
[…]
Of particular interest is the tensor product of an arbitrary GPT system and a classical GPT system, in which case there is a unique choice of a sensible tensor product corresponding to the minimal composite of GPT systems.
@@ 3.2 Operational theory associated to a GPT system
As we articulate in Section 2, proofs of (preparation) contextuality involve distinct preparation procedures, which are nevertheless operationally equivalent. Since any two distinct states of a GPT system are operationally inequivalent, we cannot identify GPT states with preparation procedures. Instead, we think of a preparation procedure as an ensemble of GPT states $\omega_{i}$, each occurring with a specified probability $\lambda_{i}$.

### v2

@@ Abstract
In this work we present a hierarchy of generalized contextuality. It refines the traditional binary distinction between contextual and noncontextual theories, and facilitates their comparison based on how contextual they are. Our approach focuses on the contextuality of prepare-and-measure scenarios, described by general probabilistic theories (GPTs). To motivate the hierarchy, we define it as the resource ordering of a novel resource theory of GPT-contextuality. The building blocks of its free operations are classical systems and univalent simulations between GPTs. These simulations preserve operational equivalences and thus cannot generate contextuality. Noncontextual theories can be recovered as least elements in the hierarchy. We then define a new contextuality monotone, called classical excess, given by the minimal error of embedding a GPT within an infinite classical system. In addition, we show that the optimal success probability in the parity oblivious multiplexing game also defines a monotone in our resource theory. Finally, we discuss whether the non-free operations can be understood as implementing information erasure and thus explaining the fine-tuning aspect of contextuality.
@@ 1 Introduction
A crucial research question in the foundations of quantum theory is to identify those features of quantum theory that constitute a true departure from the classical worldview. Addressing this question requires one to first establish a good notion of classicality, which adequately captures the classical worldview. We believe that a good notion of classicality should satisfy the following desiderata (see also [75]): (1) it endorses a principle that defines a clear boundary between aspects that pose interpretational issues and those that do not, (2) it is of broad range of applicability, (3) it is empirically testable, and (4) its violation constitutes a resource for practical applications, in particular in quantum information processing. [^1] This last desideratum is motivated by the belief that identifying the true nonclassicality of quantum theory will ultimately provide the answer to the question about the origin of the alleged quantum computational speed-up. Motivated by these desiderata, a leading notion of classicality is generalized noncontextuality [79]. [^2] In what follows we will often omit “generalized” and just talk of “(non)contextuality”. A noncontextual theory is one that is compatible with a classical realist explanation of its operational predictions —  it admits of a noncontextual ontological model. In such a model, any two experimental procedures that the theory predicts to be operationally indistinguishable also have the same ontological representation (see Section˜2 for a precise definition).
@@
A GPT system $\mathsf{B}$ is said to be at least as contextual as a GPT system $\mathsf{A}$ if there exists a univalent simulation of $\mathsf{A}$ by a composite of $\mathsf{B}$ and a classical GPT system.
It is the resource ordering of a resource theory with objects given by GPT systems and free operations given by univalent simulations with free access to classical systems. In Section˜5.1, we discuss a possible interpretation of (non-univalent) simulations as expressing a particular kind of information erasure.
@@
As is common in resource theories, the hierarchy of contextuality is not a total order — there are GPT systems such that neither is at least as contextual as the other one. It is not even a partial order, because there exist distinct GPT systems which are equivalent. For example, this is the case for all noncontextual systems, such as the classical bit and the classical trit. Therefore, the hierarchy is given by a preorder and it cannot be fully represented by a single numerical value — the “amount of contextuality" — assigned to each GPT system. However, in order to capture certain aspects of the hierarchy it is useful to define quantities which are order-preserving assignments of a number to each resource object. These are called resource monotones.
@@
Unlike Duarte and Amaral, we use the framework of GPTs. A GPT system consists of a collection of states, effects and a probability assigned to each pair of a state and an effect. While noncontextuality is traditionally phrased in terms of operational theories, it can be extended to GPTs [65] and characterized by simplex embeddability [72]. This condition states that a GPT system is noncontextual if and only if it can be embedded within a classical GPT system and thus characterizes the qualitative divide between contextual and noncontextual GPT systems. Our hierarchy of contextuality can be seen as a refinement thereof. For this purpose, we use univalent simulations between GPT systems, which are generalizations of noncontextual ontological models [64].
@@
In Section˜2, we recall the standard treatment of generalized noncontextuality in the framework of operational theories and ontological models. In order to connect it to the language of GPTs (Section˜3.1), we introduce the notion of an operational theory associated to a GPT system in Section˜3.2.Readers familiar with ontological models and GPTs may choose to skip directly to Sections˜3.3 and 3.4, where we present (univalent) simulations of GPT systems and the excess measure, respectively. We then introduce our hierarchy of contextuality in Section˜4.1 and discuss its behaviour for composite systems in Section˜4.2. We prove that the classical excess is a resource monotone (Section˜4.3) and construct a monotone from the parity-oblivious-multiplexing protocol in Section˜4.4. In Section˜5.1, we discuss a possible interpretation of the non-free operations as involving information erasure, while in Section˜5.2 we discuss the relation to previous works on contextuality and GPTs, namely [64, 28, 72, 34, 33, 78, 76, 77]. We conclude with a summary of the results and an outline of possible future research directions in Section˜6.
@@ 2 Operational theories, ontological models and contextuality
In this work we consider prepare-and-measure scenarios associated with a single system. An operational theory associated with a prepare-and-measure scenario is defined by a list of possible preparations, measurements and the probabilities $\mathcal{P}(k|P,M)$ of obtaining the outcome $k$ of the measurement $M$ given that the system is prepared in the preparation $P$. An ontological model of an operational theory provides a realist explanation of the operational predictions of the theory [43]. It does so by stipulating the existence of an ontic state space for each given system, denoted with $\Lambda$, which is mathematically represented by a (standard Borel) measurable space. Each point $\lambda\in\Lambda$ represents an ontic state that describes all the physical properties of the system. The ontological model associates each preparation $P$ in the operational theory with a conditional probability distribution $\mu_{P}(\lambda)\equiv\mu(\lambda|P)$ over ontic states. We call these epistemic states as they represent states of knowledge about the underlying ontic states. Each outcome $k$ given a measurement $M$, denoted by $[k|M]$, is associated with a conditional probability distribution $\xi(k|\lambda,M)$. The latter corresponds to the probability of obtaining outcome $k$ given that measurement $M$ is implemented on a system in the ontic state $\lambda$. An ontological model of an operational theory reproduces the predictions of the theory via the classical law of total probability,
@@
Similarly, two measurement outcomes $[k|M]$ and $k^{\prime}|M^{\prime}]$ are operationally equivalent if they give the same statistics for all possible preparations: $\mathcal{P}(k|M,P)=\mathcal{P}(k^{\prime}|M^{\prime},P)$ for all preparations $P$. In this case, we write $k|M]\simeq[k^{\prime}|M^{\prime}]$. An ontological model is measurement noncontextual if any two such measurement outcomes are represented by the same response functions:
$k|M]\simeq[k^{\prime}|M^{\prime}]\implies\xi_{[k|M]}=\xi_{[k^{\prime}|M^{\prime}]}.$ (3)
@@
We note that the notion of generalized contextuality can be also defined for transformations [79]. In this work we are only concerned with prepare-and-measure scenarios and therefore do not consider transformation contextuality.
@@ 3.1 GPT systems
Let us now discuss the mathematical description of GPT systems [60, 58, 27, 59, 53, 12] and the definition of contextuality in this context.
[…]
@@ A GPT system $\mathsf{A}$ is specified by two non-empty convex sets
$\Omega_{\mathsf{A}}\subseteq V_{\mathsf{A}},\qquad E_{\mathsf{A}}\subseteq V_{\mathsf{A}}^{*}$ (4)
that are called the state space and effect space, respectively, while their elements are states and effects of $\mathsf{A}$. We require that $E_{\mathsf{A}}$ contains the null effect $0_{\mathsf{A}}$ and the unit effect $1_{\mathsf{A}}$ satisfying
$0_{\mathsf{A}}\cdot v=0\quad\forall\,v\in V_{\mathsf{A}}\qquad\text{and}\qquad 1_{\mathsf{A}}\cdot\omega=1\quad\forall\,\omega\in\Omega_{\mathsf{A}}$ (5)
respectively. In order for the pairing to have a probabilistic interpretation, we also require $e\cdot\omega\in[0,1]$ for all effects $e$ and all states $\omega$. Finally, to avoid complications arising from degenerate systems, we require that
@@
and similarly that effects can be distinguished by states:
$e\cdot\omega=e^{\prime}\cdot\omega\quad\forall\,\omega\in\Omega_{A}\qquad\implies\qquad e=e^{\prime}.$ (8)
@@
Likewise, every classical probabilistic system is a GPT system [27, §4].
@@
$\Delta_{n}\coloneqq\mathrm{Conv}\big\{\delta_{i}\nonscript\;\,\big|\allowbreak\nonscript\;\,\mathopen{}i\in\{1,\ldots,n\}\big\},$ (11)
@@
The countable classical system $\Delta_{\mathbb{N}}$ has associated vector space $\ell^{1}$, the Banach space of sequences whose series are absolutely convergent. The state space, also denoted by $\Delta_{\mathbb{N}}$, consists of all probability measures on $\mathbb{N}$. The effect space $\Delta_{\mathbb{N}}^{*}$ consists of arbitrary sequences in $0,1]^{\mathbb{N}}$ and forms a convex subset of $\ell^{\infty}$, the space of bounded sequences, which is the dual of $\ell^{1}$.
Besides individual systems, we will occasionally also need to refer to composite ones. In general, the composite of two GPT systems is not unique — see [66, Section 5] for an in-depth discussion of tensor products of GPT systems. However, among all the possible choices, there is a ‘minimal composite’ of two GPT systems $\mathsf{A}$ and $\mathsf{B}$, whose state and effect space merely contain the separable ones. Any meaningful composite of two GPT systems necessarily contains their minimal composite as a subsystem.
@@
Given two GPT systems $\mathsf{A}=(\Omega_{\mathsf{A}},E_{\mathsf{A}},V_{\mathsf{A}})$ and $\mathsf{B}=(\Omega_{\mathsf{B}},E_{\mathsf{B}},V_{\mathsf{B}})$ we define the minimal composite system $\mathsf{A}\otimes\mathsf{B}$ to be given by the separable states and separable effects on $\mathsf{A}$ and $\mathsf{B}$. Formally $\mathsf{A}\otimes\mathsf{B}\coloneqq(\Omega_{\mathsf{A}\otimes\mathsf{B}},E_{\mathsf{A}\otimes\mathsf{B}},V_{\mathsf{A}\otimes\mathsf{B}})$ where,
@@
$\Omega_{\mathsf{A}\otimes\mathsf{B}}$ $\coloneqq\mathrm{Conv}\big\{\omega_{\mathsf{A}}\otimes\omega_{\mathsf{B}}\nonscript\;\,\big|\allowbreak\nonscript\;\,\mathopen{}\omega_{\mathsf{A}}\in\Omega_{\mathsf{A}}\text{ and }\omega_{\mathsf{B}}\in\Omega_{\mathsf{B}}\big\},$ (16)
$E_{\mathsf{A}\otimes\mathsf{B}}$ $\coloneqq\mathrm{Conv}\big\{e_{\mathsf{A}}\otimes e_{\mathsf{B}}\nonscript\;\,\big|\allowbreak\nonscript\;\,\mathopen{}e_{\mathsf{A}}\in E_{\mathsf{A}}\text{ and }e_{\mathsf{B}}\in E_{\mathsf{B}}\big\},$ (17)
Of particular interest is the tensor product of an arbitrary GPT system and a classical GPT system, in which case there is a unique choice of a sensible tensor product corresponding to the minimal composite of GPT systems (see Section˜4.2 for further discussion about this point).
@@ 3.2 Operational theory associated to a GPT system
As we articulate in Section˜2, proofs of (preparation) contextuality involve distinct preparation procedures, which are nevertheless operationally equivalent. Since any two distinct states of a GPT system are operationally inequivalent, we cannot identify GPT states with preparation procedures. Instead, we think of a preparation procedure as an ensemble of GPT states, each occurring with a specified probability. Similarly, proofs of measurement contextuality involve distinct measurement outcomes, which are operationally equivalent. Since any two distinct effects of a GPT system are operationally inequivalent, we cannot identify GPT effects with measurement outcomes. Instead, we think of a measurement procedure as a collection of GPT effects, each of which is a possible outcome of the given measurement. See [65, Section 3] for more details.
Such preparations and measurement outcomes then provide a notion of a canonical operational theory associated to a GPT system, [^4] Conversely, one can always obtain a GPT system from an operational theory. Specifically, each state is an equivalence class of preparation procedures with respect to operational equivalence, and similarly one can define the effects [44]. to which we can assign ontological models as in Section˜2. In this way, we obtain a canonical notion of an ontological model for GPT systems and a corresponding notion of noncontextuality [65, Definitions 2 and 3].
3.3 Simulations of GPT systems
We are, however, interested in comparing GPT systems and not just in distinguishing the noncontextual ones. For this purpose, we use the notion of (univalent) simulations.
Definition 6 ([64, Definition 1]).
  • a multi-valued function $\Gamma:\Omega_{\mathsf{A}}\to\Omega_{\mathsf{B}}$ called state $\epsilon$-simulation and
  • a multi-valued function $\Theta:E_{\mathsf{A}}\to E_{\mathsf{B}}$ called effect $\epsilon$-simulation,
which is empirically adequate up to an error $\epsilon\in[0,\infty)$:
$\big\lvert e\cdot\omega-f\cdot\gamma\big\rvert\leq\epsilon\qquad\forall\,\omega\in\Omega_{\mathsf{A}},\;\forall\,e\in E_{\mathsf{A}},\;\forall\,\gamma\in\Gamma(\omega),\;\forall\,f\in\Theta(e),$ (18)
preserves convex mixtures: [^5] Analogously, we require that the same holds for mixtures of effects rather than states, just as in [64].
$\lambda\Gamma(\omega)+(1-\lambda)\Gamma(\omega^{\prime})\subseteq\Gamma\bigl(\lambda\omega+(1-\lambda)\omega^{\prime}\bigr)\qquad\text{ for all }\lambda\in[0,1]\text{ and }\omega,\omega^{\prime}\in\Omega_{\mathsf{A}},$ (19)
and also preserves the null effect, i.e. we have $0_{\mathsf{B}}\in\Theta(0_{\mathsf{A}})$.
A 0-simulation is called an exact simulation. [^6] Sometimes, we refer to an exact simulation also just as a simulation for brevity. If both $\Gamma$ and $\Theta$ are single-valued, we say that the simulation is univalent.
One of the immediate consequences of the definition is that the images of states under a state simulation (and of effects under an effect simulation) are non-overlapping. Indeed, by empirical adequacy and the fact that $\Theta(e)$ is always a non-empty subset of $E_{\mathsf{B}}$, we find
$\begin{split}\gamma\in\Gamma(\omega_{1})\text{ and }\gamma\in\Gamma(\omega_{2})&\implies\forall e\in E_{\mathsf{A}}\;:\;e\cdot\omega_{1}=f\cdot\gamma=e\cdot\omega_{2}\\ &\implies\omega_{1}=\omega_{2},\end{split}$ (20)
where the second implication is by (7).